Differentiate the following function:

f(x)=\frac{ln(3x^2+5)}{5e^{2x-7}} \cdot \frac{4e^{3x}-ln(x^{2})}{(2x-x^{3})^2}
Answer #1

5e22- 4S . f(u) = ln (3x²+5) 4e3n_in(x2) 5e?-7 21-2322 = UV sery, = f(X) = u v+ uv - 0. Now. u = ln (33*+5) w sea 64_ \n(Hence from . . f(x) = UV+uv – in(3u+5) (2-x2)(124e342)-212-31?) (443x11 (92)) (2x– x3y3 + 4e3n_ in (2²) 62-2(30+5) In (3 ni